The balance of selected accounts for Huskies#1 Company at December 31, 2022,

Retained Earnings 1/1/2022 $17,200,

Dividends                               $6,000,

Sales Revenue                  $20,400

Total expenses                       $10,000

How much is the balance of retained earnings on 12/31/2022? Show all your work
Answer #1

answer:$21600

Retained earnings at 12/31/2022=Retained Earnings at 1/1/2022 + net income during the period - dividends

Retained Earnings at 1/1/2022=$17200

dividends=$6000

net income during the period=sales revenue - total expenses

sales revenue=$20400

total expenses=$10000

net income during the period=$20400 - $10000=$10400

Retained earnings at 12/31/2022=$17200 + $10400 - $6000 = $21600